OpenStudy (anonymous):

3 years ago my brother was twice as old as my sister. Now he is 12 years older than she is. How old is my sister now?

OpenStudy (lalaly):

bros age=x sister age =y (x-3)=2y x=y+12 solve now

OpenStudy (thomas5267):

TeX version of it. \[ \begin{align*} (x-3)&=2y \\ x+12&=y \end{align*} \] @lalaly \(x=y+12?\) That doesn't seems to be correct but I am not sure.

OpenStudy (anonymous):

it looks right although i'm not too sure

OpenStudy (lalaly):

hes 12 years older than his sister means his age= sisters age+12

OpenStudy (lalaly):

im sure:)

OpenStudy (anonymous):

I think that lalaly is right (: Thank you

OpenStudy (thomas5267):

Yes you are right. @lalaly

OpenStudy (lalaly):

:):)

OpenStudy (anonymous):

correct nice job